OpenStudy (anonymous):

Solve the equation by completing the square. x^2 - 2 = 9x

hero (hero):

1. Rewrite as x^2 - 9x = 2 Let b = -9 2. add (b/2)^2 to both sides: x^2 -9x + (-9/2)^2 = -9 + (-9/2)^2 x^2 -9x + 81/4 = -9 + 81/4 3. Re-write the left side as a binomial square: (x - 9/2)^2 = -9 + 81/4 4. Simplify the right side: (x - 9/2)^2 = -36/4 + 81/4 (x - 9/2)^2 = (81-36)/4 (x - 9/2)^2 = (45)/4 5. Square root both sides to get x - 9/2 = + or - sqrt(45/4) 6. Add 9/2 to both sides x = + or - sqrt(45/4) + 9/2

OpenStudy (anonymous):

thanks man!